logx+4 125=3 .. helpp
Answer by RAY100(1637)

log x +4.125 =3,,,,,,pls check problem
.
log x = 3-4.125 = -1.125
.
x= .075
.
check,,,log .075 +4.125 =3
.
-1.125 +4.125 =3,,,,,,ok
.
log functions become undefined as x approaches 0.
.
we assume the decimal,,4.125, which allows for a reasonable answer